Zn(CN),: 4.95 x 10-4 g/L Express your answer using three significant figures. IVO AQ R o E ?
Answer #1

Concentration of ancem2 = (4.959) x(11704449 mal =1) x 10-4 molar mars of 2n Cenda = 4021810-70 mol/L = 117 444 g/mol aneen)2